ID |
Date |
Author |
Type |
Category |
Subject |
2668
|
Thu Mar 11 17:51:04 2010 |
Koji | Update | SUS | Recent status of SOSs |
Jenne, Koji
Recent status of SOSs:
We completed one of the suspension (ITMY).
ITMX: 6 Magnets, standoffs, and guide rod glued / balance to be confirmed / needs to be baked
ITMY: 6 Magnets, standoffs, and guide rod glued / balance confirmed / needs to be baked
SRM: 6 Magnets, one standoff, and guide rod glued, / waiting for the release from the gluing fixture.
PRM: one standoff, and guide rod glued / waiting for the magnet gluing.
We think we solved all the problems for hanging the suspensions.
--- Magnet gluing fixture ---
- There is the two kinds of fixtures. Neither does work propery in the original form!
- The height of the side magnets should be finely adjusted by changing the teflon sheets beneath the optics in the fixture.
- Be aware of the polarity of the fixture in terms of the side magnets
- Wrongly glued magnets (and others) can be removed by a razor blade with some amount of acetone.
- The pickle picker frequently knocks the magnets down during the release. Don't s be down in the dumps too much.
--- Suspending the mirror ---
- The wire winches must be carefully attached to the suspension tower such that the wires are not streached during fastening the clamps.
- There are a couple variations of the drawings for SOS. The one we have has #4-40 for the earthquake stops at the bottom.
Zach and Mott made the EQ stops with the right size.
|
2667
|
Thu Mar 11 15:24:51 2010 |
steve | Metaphysics | Environment | IFO was well protected |
Quote: |
There is a planned power outage tomorrow, Saturday from 7am till midnight.
I vented all annulies and switched to ALL OFF configuration. The small region of the RGA is still under vacuum.
The vac-rack: gauges, c1vac1 and UPS turned off.
|
It turns out that we perfecly timed the big one
In the process of finding the signal of the big chilean earthquake I just realized that we were all off |
2666
|
Wed Mar 10 15:02:58 2010 |
rana | Update | PEM | Styrofoam Cooler on the Seismos |
This is the spectra and coherence from a quiet time last night. I've lowered the Guralp cal by a factor of 2 to account for the fact that the gain in the breakout box is actually 20 and not 10 as I previously said.
The AD620 stage in the front part has a gain of 10 and then there's a single-to-differential stage in the output which gives us a gain of 2. The DTT cal in counts is now 3.8e-9 (m/s)/count.
The second plot shows the Guralp and Ranger signals at the ADC input (converted from counts to Volts for usefulness). The thick grey line is the expected noise of the Guralp breakout box
(mainly the AD620) propagated to the ADC (via multiplication by 2). It looks like the preamp board should not be a problem as long as we can reach the AD620 limit.
So the excess noise in the Guralp is not the fault of the preamp, but more likely the mounting and insulation of the seismometers. |
Attachment 1: Seismic-Ref_100310.png
|
|
Attachment 2: Untitled.gif
|
|
2665
|
Tue Mar 9 12:06:53 2010 |
rana, Jenne | Update | PEM | Styrofoam Cooler on the Seismos |
Looks like the GUR2_X signal is bad. Jenne says that we need to center it mechanically before the signals will become useful again. Maybe Steve will do this - instructions are in the manuel ? |
2664
|
Tue Mar 9 09:32:31 2010 |
Koji | Summary | General | Wideband measurement of Fast PZT response |
I have measured a wideband response of the fast PZT in the LWE NPRO 700mW in the Alberto's setup.
This is a basic measurement to determine how much phase modulation we can obtain by actuating the fast PZT,
primarily for the green locking experiment.
RESULT
- Above 200kHz, there are many resonances that screws up the phase.
- Modulation of 0.1rad can be easily obtained even at 10MHz if the modulation frequency is scanned.
- Change of the laser frequency in DC was observed depending on the modulation frequency.
i.e. At the resonance the laser frequency escaped from the RF spectrum analyzer.
This may induced by the heat dissipation in the PZT causing the temperature change of the crystal.
- Some concerns: Is there any undesired AM by the PZT modulation?
---
METHOD
1. Locked the PLL of for the PSL-NPRO beating at 20MHz.
2. Added the modulation signal to the NPRO PZT input.
I used the output of the network analyzer sweeping from 100kHz to 1MHz.
3. Measured the transfer function from the modulation input to the PLL error signal.
The PLL error is sensitive to the phase fluctuation of the laser. Found that the first resonance is at 200kHz.
The TF is not valid below 3kHz where the PLL suppresses the modulation.
4. Single frequency modulation: Disconnected the PLL setup.
Plug Marconi into the fast PZT input and modulate it at various frequencies.
Observing with the RF spectrum analyzer, I could see strong modulation below 1MHz.
It turned out later that the TF measurement missed the narrow peaks of the resonances due to the poor freq resolution.
Also the modulation depth varies frequency by frequency because of the resonances.
Scanned the frequency to have local maximum of the modulation depth. Adjusted the
modulation amplitude such that the carrier is suppressed (J0(m)=0 i.e. m~2.4). As I could not obtain
the carrier suppression at above 1MHz, the height of the carrier and the sidebands were measured.
The modulation frequency was swept from 100kHz to 10MHz.
5. Calibration. The TF measured has been calibrated using the modulation depth obtained at 100Hz,
where the resonance does not affect the response yet.
The responce of the PZT was ~10MHz/V below 30kHz. Looks not so strange although this valure is
little bit high from the spec (2MHz/V), and still higher than my previous experience at TAMA (5MHz/V).
Note that this calibration does not effect to the modulation depth of the single freq measurement as they are independent. |
Attachment 1: PZT_response.png
|
|
2663
|
Tue Mar 9 09:04:20 2010 |
steve | Update | PEM | keep vacuum chamber closed |
They are sandblasting at CES: our particle counts are very high. DO NOT OPEN CHAMBER! |
Attachment 1: sandblasting.jpg
|
|
2662
|
Mon Mar 8 18:15:27 2010 |
rana | Update | PEM | Styrofoam Cooler on the Seismos |
I put Jenne's cooler over the seismometers. Kiwamu put the copper foil wrapped lead brick on top of the cooler to hold it down. I also put another (unwrapped) lead brick on top of the Guralp cables outside of the cooler. Frank gave me a knife with which I cut a little escape hole in the bottom of the cooler lip for the Guralp cables to sneak out of. |
Attachment 1: SeismometersHuddled.jpg
|
|
2661
|
Sun Mar 7 23:05:39 2010 |
rana | Update | PEM | Seismic witnesses near MC1 tank moved |
Quote: |
Since we're going to open the MC1 tank tomorrow, I've moved the MC1 accelerometers and the Guralp over to underneath MC2 for the vent. I'll reconnect them later.
|
I've put both Guralps next to the Ranger and connected them to the breakout box. The data is now good.
I found that the Ranger was not centered and so it was stuck (someone kicked it in the last 2 weeks apparently). I recentered the mass according to the procedure in the manual. Its now moving freely.
In order to do a better huddle test, I increased the gain of the Ranger's SR560 preamp to 100 from 10 and put it on the low noise setting. I also enabled a 2x lowpass at 3 kHz for no good reason.
I couldn't find what the actual value of the gain of the Guralp breakout box is, but I assume its 10. With this assumption the calibrations are this:
Guralp: 800 V/(m/s) * 10 (V/V) * 16384 cts/V => 7.63e-9 (m/s)/count (0.03 - 40 Hz)
Ranger: 345 V/(m/s) * 100 (V/V) * 16384 cts/V => 1.77e-9 (m/s)/count (above 1Hz)
To account for the fact that I am not damping the Ranger with an external damping resistor, I have changed the calibration poles and zeros: in DTT we now use 2 poles @ 0 Hz and a complex pair at 1 Hz:
G = 1.77e-9
Poles = 0, 0
Zeros = 0.15 0.9887
I think that the Guralp gain is too high by a factor of 2. To really do this right, we should attach a known voltage to the input pins of the Guralp breakout and then read off the amount of counts. |
Attachment 1: seis.png
|
|
2660
|
Sun Mar 7 07:01:21 2010 |
rana | Update | WienerFiltering | Guralp Huddle Test software |
We need to do a new huddle test of the Guralps for the Wiener filtering paper. The last test had miserable results.
I tried to use recent data to do this, but it looks like we forgot to turn the Guralp box back on after the power outage or that they're far off center.
So instead I got data from after the previous power outage recovery.
I tried to use our usual Wiener filter method to subtract Guralp1-Z from Guralp2-Z, but that didn't work so well. It was very sensitive to the pre-weighting.
Instead I used the new .m file that Dmass wrote for subtracting the phase noise from his doubling noise MZ. That worked very well. It does all of the subtraction in the frequency domain and so doesn't have to worry about making a stable or causal filter. As you can see, it beats our weighted Wiener filter at all frequencies.

The attached plot shows the Guralp spectra (red & green), the residual using time-domain Wiener filtering (black) and the Dmass f-domain code (yellow).
As soon as Jenne brings in her beer cooler, we're ready to redo the Huddle Test.
|
2659
|
Fri Mar 5 18:04:56 2010 |
rana | Update | SUS | used OSEMs are magnetic |
The OSEM LEDs and PDs from Honeywell have always had some ferromagnetic material in them. These are the same OSEMs we had since 2000.
You must be thinking of the really old 20th century plastic OSEMs. |
2658
|
Fri Mar 5 11:21:18 2010 |
steve | Update | SUS | used OSEMs are magnetic |
Quote: |
Jenne and Koji
We successfully hung ITMX on the SOS. Side magnet is ~2mm off from the center of the OSEM. ITMX aligned using the QPD. The OSEMs changes the alignment. It looks that something magnetic is inside the OSEM PD or LED.
Reguled ITMY side magnet.
Cleaned up the lab for the safety inspection.
|
The brand new OSEM LED and PD can be picked up with a weak magnet. These ferrous metals of LEDs and PDs will be magnetized by sitting in the sus next to the
magnets for years. I hanged optics with new OSEMs and never saw this effect before.
We have to demagnetize them. |
2657
|
Thu Mar 4 22:07:21 2010 |
rana | Update | 40m Upgrading | 11MHz PD not yet designed |
Use 10 Ohms for the resistance - I have never seen a diode with 25 Ohms.
p.s. PDFs can be joined together using the joinPDF command or a few command line options of 'gs'. |
2656
|
Thu Mar 4 19:53:56 2010 |
Alberto | Update | 40m Upgrading | 11MHz PD designed adjusted for diode's resistance; 55 MHz RFPD designed |
I read a few datasheets of the C30642GH photodiode that we're going to use for the 11 and 55 MHz. Considering the values listed for the resistance and the capacitance in what they define "typical conditions" (that is, specific values of bias voltage and DC photocurrent) I fixed Rd=25Ohms and Cd=175pF.
Then I picked the tunable components in the circuit so that we could adjust for the variability of those parameters.
Finally with LISO I simulated transfer functions and noise curves for both the 11 and the 55MHz photodiodes.
I'm attaching the results and the LISO source files.
|
Attachment 1: rfpd55_Noise.pdf
|
|
Attachment 2: rfpd55_TF.pdf
|
|
Attachment 3: rfpd11_v2_TF.pdf
|
|
Attachment 4: rfpd11_v2_Noise.pdf
|
|
Attachment 5: rfpd11_v2.fil
|
Attachment 6: rfpd55.fil
|
2655
|
Thu Mar 4 08:43:35 2010 |
Alberto | Update | 40m Upgrading | 11 MHz RFPD elctronics |
Quote: |
** Please add LISO file w/ component values.
|
oops, forgotten the third attachment...
here it is |
Attachment 1: rfpd11_v2.fil
|
# Resonant RF diode front end
#
# gnd
# |
# Cw2
# |
# n23
# |
# Lw2
# |
... 60 more lines ...
|
2654
|
Thu Mar 4 02:25:14 2010 |
Jenne | Update | COC | Further details on the magnet story, and SRM guiderod glued |
[Koji, Jenne]
First, the easy story: SRM got it's guiderod & standoff glued on this evening. It will be ready for magnets (assuming everything is sorted out....see below) as early as tomorrow. We can also begin to glue PRM guiderods as early as tomorrow.
The magnet story is not as short.....
Problem: ITMX and ITMY's side magnets are not glued in the correct places along the z-axis of the optic (z-axis as in beam propagation direction).
ITMX (as reported the other day) has the side magnet placement off by ~2mm. ITMX side was glued using the magnet fixture from MIT and the teflon pads that Kiwamu and I improvised.
It was determined that the improvised teflon pads were too thin (maybe about 1m thick), so I took those out, and replaced them with the teflon pads stolen from the 40m's magnet gluing fixture. (The teflon pad from the MIT fixture and the ones from the MIT fixture are the same within my measuring ability using a flat surface and feeling for a step between them. I haven't yet measured with calipers the MIT pad thickness). The pads from the 40m fixture, which were used in the MIT fixture to glue ITMY side last night were measured to be ~1.7mm thick.
Today when Koji hung ITMY, he discovered that the side magnet is off by ~1mm. This improvement is consistent with the switching of the teflon pads to the ones from the 40m fixture.
We compared the 40m fixture with the one from MIT, and it looks like the distance from the edge of where the optic should sit to the center of the hole for the side magnet is different by ~1.1mm. This explains the remaining ~1mm that ITMY is off by.
We should put the teflon pads back into the 40m fixture, and only use that one from now on, unless we find an easy way to make thicker teflon pads for the fixture we received from MIT. (The pads that are in there are about the maximum thickness that will fit). I'm going to use my thickness measurements of SRM (taken in the process of gluing the guiderods) to see what thickness of pads / what fixture we want to actually use, but I'm sure that the fixture we found in the 40m is correct. We can't use this fixture however, until we get some clean 1/4-28 screws. I've emailed Steve and Bob, so hopefully they'll have something for us by ~lunchtime tomorrow.
The ITMX side magnet is so far off in the Z-direction that we'll have to remove it and reglue it in the correct position in order for the shadow sensor to do anything. For ITMY, we'll check it out tomorrow, whether the magnet is in the LED beam at all or not. If it's not blocking the LED beam enough, we'll have to remove and reglue it too.
Why someone made 2 almost identical fixtures, with a 1mm height difference and different threads for the set screws, I don't know. But I don't think whoever that person was can be my friend this week. |
2653
|
Wed Mar 3 18:32:25 2010 |
Alberto | Update | 40m Upgrading | 11 MHz RFPD elctronics |
** Please add LISO file w/ component values.
I designed the circuit for one of the 11 MHz photodiodes that we're going to install in the 40m Upgrade.
This is a simple representation of the schematic:
gnd # | # Cw2 # | # n23 # | # Lw2 # | # n22 # | # Rw2 # | |\ # n2- - - C2 - n3 - - - - | \ # | | | | |4106>-- n5 - Rs -- no # iinput Rd L1 L2 R24 n6- | / | | # nin - | | | | | |/ | Rload # Cd n7 R22 gnd | | | # | | | | - - - R8 - - gnd # gnd R1 gnd R7 # | | # gnd gnd # # #
I chose the values of the components in a realistic way, that is using part available from Coilcraft or Digikey.
Using LISO I simulated the Tranfer Function and the noise of the circuit.
I'm attaching the results.
I'll post the 55MHz rfpd later. |
Attachment 1: rfpd11_v2_TF.pdf
|
|
Attachment 2: rfpd11_v2_Noise.pdf
|
|
2652
|
Wed Mar 3 02:57:29 2010 |
Jenne | Update | SUS | ITMX hung |
Quote: |
Jenne and Koji
We successfully hung ITMX on the SOS. Side magnet is ~2mm off from the center of the OSEM.
Reguled ITMY side magnet.
Cleaned up the lab for the safety inspection.
|
Some details on the side magnet situation from today:
To glue the magnets+dumbbells to the optics, we use the magnet-dumbbell gluing fixture. This fixture is supposed to have teflon 'pads' for the optic to sit on while you align it in the fixture, however the fixture which we received from MIT (it's Betsy's....but it came via MIT) only had one of the 4 teflon pads.
Kiwamu and I decided (last week, when we first glued ITMX's magnets) that it would be bad news to let the AR face of the optic sit on bare metal, so we fashioned up some teflon pads using stock in a cabinet down the Yarm. We were focused on thinking about the face magnets, and didn't think about how the thickness of the teflon affected the placement of the side magnet. We chose some teflon that was too thin by ~1mm, so the optic sat too low in the fixture, resulting in the side magnet being glued too close to the HR side of the optic (this is all along the Z - axis, where Z is the direction of beam propagation).
Why it ended up being 2mm off instead of only 1mm I don't really have an explanation for, other than perhaps tightening the set screws to hold the optic (by the barrel) in the fixture pushes the optic up. I observed this happening when I didn't put any effort into keeping the optic flat on the teflon pads, but I thought that I made sure the optic was seated nicely in the fixture before starting to glue. When I glued the new ITMY side magnet tonight I tried to make sure that the optic was seated nicely in the fixture. We'll see what happens.
Before gluing the new ITMY side magnet (and now it's set for all future magnet gluings....), I found 4 teflon pads of all the correct thickness. It turns out that we have a magnet gluing fixture of our own, which I found in the cabinets in the clean room. This fixture had all 4 teflon pads, so I stole them and put them into the one that we're using for this round of upgrade / suspension hangings. The height of all future side magnets should be correct. The thickness of the pads in the 'spare' fixture matched the one which came with the fixture from MIT as closely as I could feel by putting them on the same flat surface next to each other and feeling if there was a step.
A side note about this magnet gluing fixture that I found: It has the word "TOP" etched into it, to prevent exactly my problem with the ITMY side magnets in the first place. Unfortunately the threads for the set screws which hold the optic are shot (or something is funny with them), so we can't just use this fixture.
Gluing notes regarding the standoffs and guiderods:
There's more glue than I'd like on the guiderods / standoff for ITMX. The glue was starting to get a little tacky when I glued the standoff in place after we balanced the optic, so it was hard to get it in the right place. I'm confident we have a good epoxy contact, and we don't have much glue that I think it'll be a big problem. Certainly I'll be a lot better at manuvering my glue-stick a.k.a skinny piece of wire around the suspension tower to get to the standoff for the rest of the optics that we're hanging, and I won't have glued something like ITMY side magnet immediately beforehand, which took enough time that the glue started to get tacky (not very tacky, just barely noticeably tacky).
I'd say that most gluing activities should be completed within ~10-15min of mixing the glue, after spending ~2min stirring to make sure it's nice and uniform. It doesn't dry fast enough to be a huge rush, but you should get right on the gluing once the epoxy has been mixed. |
2651
|
Tue Mar 2 23:11:43 2010 |
Koji | Update | SUS | ITMX hung |
Jenne and Koji
We successfully hung ITMX on the SOS. Side magnet is ~2mm off from the center of the OSEM. ITMX aligned using the QPD. The OSEMs changes the alignment. It looks that something magnetic is inside the OSEM PD or LED.
Reguled ITMY side magnet.
Cleaned up the lab for the safety inspection. |
2650
|
Tue Mar 2 12:20:54 2010 |
kiwamu | Update | PSL | stray beam |
In order to block stray beams, I have put some beam dumps and razor blades on the PSL table.
There were three undesired spots in total. I found two spots on the south side door of the PSL room, close to Mach-Zehnder.
Another spots was on the middle of the north door. Now they all are blocked successfully. |
2649
|
Mon Mar 1 22:38:12 2010 |
rana | Update | Computers | RC sensitivity to RIN |
The overnight triangle wave I ran on the AOM drive turns out to have produced no signal in the FAST feedback to the PZT.
The input power to the cavity was ~10 mW (I'm totally guessing). The peak-peak amplitude of the triangle wave was 50% of the total power.
The spectral density of the fast signal at the fundamental frequency (~7.9 mHz) is ~0.08 V/rHz. The FAST calibration is ~5 MHz/V. So, since we
see no signal, we can place an upper limit on the amount of frequency shift = (5 MHz/V) * (0.08 V/rHz) * sqrt(0.0001 Hz) = 4 kHz.
Roughly this means that the RIN -> Hz coefficient must be less than 4 kHz / 5 mW or ~ 1 Hz/uW.
For comparison, the paper on reference cavities by the Hansch group lists a coefficient of ~50 Hz/uW. However, they have a finesse of 400000
while we only have a finesse of 8000-10000. So our null result means that our RC mirrors' absorption is perhaps less than theirs. Another possibility
is that their coating design has a higher thermo-optic coefficient. This is possible, since they probably have much lower transmission mirrors. It would be
interesting to know how the DC thermo-optic coefficient scales with transmission for the standard HR coating designs. |
Attachment 1: Untitled.png
|
|
2648
|
Mon Mar 1 11:21:50 2010 |
steve | Configuration | VAC | vacuum contoller is back to normal |
Quote: |
There is a planned power outage tomorrow, Saturday from 7am till midnight.
I vented all annulies and switched to ALL OFF configuration. The small region of the RGA is still under vacuum.
The vac-rack: gauges, c1vac1 and UPS turned off.
|
Vac- rack is powered back up. UPS first, than all other power switches from top to bottom of the rack, except Maglev
Manually started one by one TP2 and TP3 to accelerate to 50 KRPM
Brought up vac.control screen on lap-top at /cvs/cds/caltech/medm/c0/ve/VacControl_BAK.adl
V5 and VM3 were opened so TP3 can pump on the RGA
V4 was opened so TP2 can pump on the Maglev-TP1. The Maglev power was turned on and started acceleration.
The vac control screen positions indicators were checked for true position and annulies vent valves were opened.
RGA manual on/off switch was turned at the top of the RGA-head. Ubuntu copmuter was started at cc4 1.1e-6 Torr
The RGA communication was started with: ssh c0rga from control room
The rga-script was started ./RGAset.py This script turns on the filament, rs-232 and scan parameters etc
Vac -configuration: IFO-P1 at atm, RGA is pumped and running in background mode, all annulos at atm
|
2647
|
Mon Mar 1 08:49:37 2010 |
steve | Bureaucracy | SAFETY | safety audit |
The 40m safety audit will be at Wednesday afternoon, March 3
Please participate getting our lab to inspection grade safety level. |
2646
|
Sun Feb 28 23:47:52 2010 |
rana | Update | Computers | rosalba |
Since Rosalba wanted to update ~500 packages, I let it do it. This, of course, stopped the X server from running. I downloaded and installed the newest Nvidia driver and its mostly OK.
The main problem with the auto-update on our workstations is that we've updated some packages by hand; i.e. not using the standard CentOS yum. So that means that the auto-update doesn't work right. From now on, if you want to install a fancier package than what CentOS distributes, you should commit to handle the system maintenance for these workstations for the future. Its not that we can't have new programs, we just have to pay the price.
At 23:45 PST, I also started a slow triangle wave on the AOM drive amplitude. This is to see if there's a response in the FSS-FAST which might imply a coupling from intensity noise to frequency noise via absorbed power and the dn/dT effect in the coatings.
Its a 93 second period triangle modulating the RC power from 100% down to 50%. |
2645
|
Sun Feb 28 16:45:05 2010 |
rana | Summary | General | Power ON Recovery |
- Turned ON the RAID above linux1.
- Hooked up a monitor and keyboard and then turned ON linux1.
- After linux1 booted, turned ON nodus - then restarted apache and elog on it using the wiki instructions.
- Turned on all of the control room workstatiions, tuned Pandora to Johnny Cash, started the auto package updater on Rosalba (517 packages).
- Started the startStrip script on op540m.
- turned on RAID for frames - wait for it to say 'SATA', then turn on daqctrl and then fb40m and then daqawg and then dcuepics
- turned on all the crates for FEs, Sorensens, Kepcos for LSC, op340m, mafalda was already on
- fb40m again doesn't mount the RAID again!
- I turned on fb40m2 and that fixes the problem. The fb40m /etc/vfstab points to 198.168.1.2, not the JetStor IP address.
- I plugged in the Video Switch - its power cord was disconnected.
- FEs still timing out saying 'no response from EPICS', but Alberto is now here.
Sun Feb 28 18:23:09 2010
Hi. This is Alberto. Its Sun Feb 28 19:23:09 2010
- Turned on c1dcuepics, c0daqctrl and c0daqawg. c0daqawg had a "bad"status on the daqdetail medm screen. The FEs still don't come up.
- Rebooted c1dcuepics and power cycled c0daqctrl and c0daqawg. The problem is still there.
- Turned on c1omc. Problem solved.
- Rebooted c1dcuepics and power cycled c0daqctrl and c0daqawg. c0daqawg now good. The FE are coming up.
- Plugged in the laser for ETMY's oplev
- Turned on the laser of ETMX's oplev from its key.
Monday, March 1, 9:00 2010 Steve turns on PSL-REF cavity ion pump HV at 1Y1 |
2644
|
Fri Feb 26 15:32:13 2010 |
steve | Configuration | VAC | preparation for power outage: vacuum all off |
There is a planned power outage tomorrow, Saturday from 7am till midnight.
I vented all annulies and switched to ALL OFF configuration. The small region of the RGA is still under vacuum.
The vac-rack: gauges, c1vac1 and UPS turned off. |
Attachment 1: ventd3.jpg
|
|
2643
|
Fri Feb 26 11:48:36 2010 |
Koji | Update | General | MC incident beam shift |
Last night I worked on the MC incident beam such that we can hit the center of the MC mirrors.
Steve and I checked the incident beam on MC1. We found the beam is ~5mm south.
This was not too critical but it is better to be realigned. I moved the steering mirror on the OMC
table (in vac). We kept the MC resonated. After the maximization of the resonance, I realigned the
MC1 and MC3 such that the resonance in dominated by TEM00.
Jenne, Kiwamu, and I then closed the light door on to the OMC/IMC.
I will make more detailed entry with photos in order to explain what and how I did. |
2642
|
Fri Feb 26 01:00:07 2010 |
Jenne | Update | COC | Suspension Progress |
This is going to be a laundry list of the mile markers achieved so far:
* Guiderod and wire standoff glued to each ITMX and ITMY
* Magnets glued to dumbbells (4 sets done now). ITMX has 244 +- 3 Gauss, ITMY has 255 +- 3 Gauss. The 2 sets for SRM and PRM are 255 +- 3 G and 264 +- 3 G. I don't know which set will go with which optic yet.
* Magnets glued to ITMX. There were some complications removing the optic from the magnet gluing fixture. The way the optic is left with the glue to dry overnight is with "pickle picker" type grippers holding the magnets to the optic. After the epoxy had cured, Kiwamu and I took the grippers off, in preparation to remove the optic from the fixture. The side magnet (thankfully the side where we won't have an OSEM) and dumbbell assembly snapped off. Also, on the UL magnet, the magnet came off of the dumbbell (the dumbbell was still glued to the glass). We left the optic in the fixture (to maintain the original alignment), and used one of the grippers to glue the magnet back to the UL dumbbell. The gripper in the fixture has very little slop in where it places the magnet/dumbbell, so the magnet was reglued with very good axial alignment. Since after the side magnet+dumbbell came off the glass, the 2 broke apart, we did not glue them back on to the optic. They were reattached, so that we can in the future put the extra side magnet on, but I don't think that will be necessary, since we already know which side the OSEM will be on.
* Magnets glued to ITMY. This happened today, so it's drying overnight. Hopefully the grippers won't be sticky and jerky like last time when we were removing them from the fixture, so hopefully we won't lose any magnets when I take the optic out of the fixture.
* ITMX has been placed in its suspension cage. The first step, before getting out the wire, is to set the optic on the bottom EQ stops, and get the correct height and get the optic leveled, to make things easier once the wire is in place. Koji and I did this step, and then we clamped all of the EQ stops in place to leave it for the night.
* The HeNe laser has been leveled, to a beam height of 5.5inches, in preparation for the final leveling of the optics, beginning tomorrow. The QPD with the XY decoder is also in place at the 5.5 inch height for the op lev readout. The game plan is to leave this set up for the entire time that we're hanging optics. This is kind of a pain to set up, but now that it's there, it can stay out of the way huddled on the side of the flow bench table, ready for whenever we get the ETMs in, and the recoated PRM.
* Koji and Steve got the ITMX OSEMs from in the vacuum, and they're ready for the hanging and balancing of the optic tomorrow. Also, they got out the satellite box, and ran the crazy-long cable to control the OSEMs while they're on the flow bench in the clean room.
Koji and I discovered a problem with the small EQ stops, which will be used in all of the SOS suspensions for the bottom EQ stops. They're too big. :( The original document (D970312-A-D) describing the size for these screws was drawn in 1997, and it calls for 4-40 screws. The updated drawing, from 2000 (D970312-B-D) calls for 6-32 screws. I naively trusted that updated meant updated, and ordered and prepared 6-32 screws for the bottom EQ stops for all of the SOSes. Unfortunately, the suspension towers that we have are tapped for 4-40. Thumbs down to that. We have a bunch of vented 4-40 screws in the clean room cabinets, which I can drill, and have Bob rebake, so that Zach and Mott can make viton inserts for them, but that will be a future enhancement. For tonight, Koji and I put in bare vented 4-40 screws from the clean room supply of pre-baked screws. This is consistent with the optics in our chambers having bare screws for the bottom EQ stops, although it might be nicer to have cushy viton for emergencies when the wire might snap. The real moral of this story is: don't trust the drawings. They're good for guidelines, but I should have confirmed that everything fit and was the correct size. |
2641
|
Thu Feb 25 19:59:50 2010 |
Koji | Configuration | SUS | ITMX OSEMs |
Koji, Steve
ITMX OSEM CONFIGURATION
|
Attachment 1: osem_config_ITMX.png
|
|
2640
|
Thu Feb 25 15:49:05 2010 |
Alberto | AoG | CDS | New IO Chassis for the new CDS |
Yesterday Kiwamu and I went to Downs to take all the available parts of the IO chassis that Gary and I had put together over there.
We've got only 3 of the 5 that we need for the Upgrade. The other 2 are currently being used for some other purpose in Downs labs.
I'm not sure about what each chassis has supposed to contain. They all also look different from each other.
Anyway, it looks like there should be a sort of motherboard and an IO Chassis Interface Board (DCC# D0902029) in each of them. The IO Chassis Interface Board is just a board with a bunch of PCI slots.
This is what the 3 chassis that we've got yesterday have:
Chassis 1
- 1 very big "motherboard"
- power supply
Chassis 2
- small motherboard
- IO Interface Board (DCC# D0902029)
- power supply
Chassis n.3
- "Dolpjin" motherboard
- IO Interface Board
- power supply
Apparently 2 of these 3 chassis are still missing their IO interface boards,
Also all chassis are still missing all the connections to powering, fans, LEDs, power and reset buttons. It's not clear how these connections should be. Gary didn't know it either. |
2639
|
Thu Feb 25 11:21:06 2010 |
Koji | Update | General | Tanks opened |
[Steve, Bob, Joe, Zach, Alberto, Kiwamu, Koji]
We opened the OMC-IMC access connector, ITMX North door, and ITMY West door.
We worked from 9:30-11:00.
The work was quite smooth thanks to the nice preparation of Steve as usual.
Thank the team for the great work!
|
2638
|
Wed Feb 24 16:11:15 2010 |
Jenne | Update | PEM | Seismic witnesses near MC1 tank moved |
Since we're going to open the MC1 tank tomorrow, I've moved the MC1 accelerometers and the Guralp over to underneath MC2 for the vent. I'll reconnect them later. |
2637
|
Wed Feb 24 12:08:31 2010 |
Koji | Update | Computers | RFM goes red -> recovered by the nuclear option |
Most of the RFM went red this morning. I took the nuclear option and it seemed to be recovered. |
2636
|
Tue Feb 23 19:07:43 2010 |
kiwamu and steve | Configuration | VAC | vent is completed |
Quote: |
Kiwamu and Steve have started venting the 40m vacuum envelope.
Preparation:
centered oplevs at resonating cavities,
ITM references were set by green pointer from the ends by Koji,
closed PSL shutter and placed manual block into beam path,
checked jamnuts in locked positions on bellows,
turned HV off at PZT-Jena "steering mirror" power supply and OMC HV ps
checked particle counts,
switched oplev servos off,
set up N2 cylinder to start vent from 1e-6 Torr to 25 Torr,
have ~ 6 cylinders of instrument grade compressed air to bring envelope from 25 Torr to 760 Torr
All three cranes were wiped off today.
|
Kiwamu has completed the vent. |
Attachment 1: vent2010feb23.jpg
|
|
2635
|
Tue Feb 23 19:00:45 2010 |
kiwamu | Configuration | VAC | vent finished |
The vent has been finished.
Now the pressure inside the chamber is 760 torr, and it's getting equilibrium with the atmospheric pressure.
Therefore we are ready and can open the door of the chamber tomorrow. |
2634
|
Tue Feb 23 16:42:02 2010 |
rana | Configuration | Computer Scripts / Programs | SVN restarted on NODUS |
I ran the start Apache script as described by Yoichi in the WIki. SVN back up. |
2633
|
Tue Feb 23 16:20:44 2010 |
steve | Configuration | VAC | outgassing + leak |
Quote: |
As preparation for the upcoming planned power outage we turned turbos, RGA off and closed valves.
IFO chamber is not pumped now. Small leaks and out gassing will push the pressure up slowly. At 3 mTorr of P1 the PSL output shutter
will be closed by the interlock.
It is OK to use light in the IFO up to this point.
|
Our vacuum was not pumped for 4 days. The computers were not up when we started pumping again. The manual reading on P1 was 15 mTorr.
This means that our outgassing plus leakrate is ~3.8 mTorr /day
3-5 mTorr / day is normal |
Attachment 1: outgleak.jpg
|
|
2632
|
Tue Feb 23 14:56:24 2010 |
steve | Update | General | ETMX optical table mass |
Our janitor dropped one 48" long fluorescent tube on the top cover of ETMX-isct. This accident made glasses fly all over the place.
He cleaned up nicely, but please beware of small glass pieces around ETMX chamber.
We did not clean up on the table in order to reserve oplev and green ITMX pointing alignment.
Plot below shows that the alignment was not effected.
|
Attachment 1: etmxmass.jpg
|
|
2631
|
Tue Feb 23 13:37:04 2010 |
kiwamu and steve | Configuration | VAC | venting the 40m vac envelope |
Kiwamu and Steve have started venting the 40m vacuum envelope.
Preparation:
centered oplevs at resonating cavities,
ITM references were set by green pointer from the ends by Koji,
closed PSL shutter and placed manual block into beam path,
checked jamnuts in locked positions on bellows,
turned HV off at PZT-Jena "steering mirror" power supply and OMC HV ps
checked particle counts,
switched oplev servos off,
set up N2 cylinder to start vent from 1e-6 Torr to 25 Torr,
have ~ 6 cylinders of instrument grade compressed air to bring envelope from 25 Torr to 760 Torr
All three cranes were wiped off today.
|
2630
|
Tue Feb 23 06:47:57 2010 |
Koji | Update | General | IFO situations / low power MC lock |
Work on 22nd Monday:
[MC recovery]
- Tried to lock MC after the computer recovery by Joe.
- A lot of higher modes. I can touch the input periscope or the MC mirrors.
- First tried to align the MC mirrors. MC1 was aligned against the MC REFL PD. MC2/3 was aligned to maximize the transmitted power.
- After the alignment, I got the MC Trans Sum ~8V. Also I saw the flashing of the arm cavities. I decided to take this alignment although the beam looks little bit clipped by the faraday.
[IFO alignment recovery]
- Aligned the arms for TEM00 manually.
- Arm alignment script seems not working now. This could come from the move of the end QPDs
- PRMI/DRMI were aligned. All alignment values saved.
[Low power MC]
[Optical config]
- I fixed the MCT CCD camera. It is quite useful to align the MC.
- Inserted HWP+Cube PBS+HWP combo in the MC incident path.
- First HWP and PBS adjust the light power. The second HWP is fixed at 342deg such that it restores the poralization to S.
- The incident power was measured by the SCIENTECH power meter. Offset of 3mW was subtracted in the table below.
HWP1 angle |
P_MC_incident |
comment |
126deg |
1.03W |
Max |
100 |
0.39 |
|
90 |
0.098 |
Low power max |
85 |
0.021 |
Low power nominal |
- HWP1 85deg is the nominal.
- I needed to touch the steering mirror (indicated by the picture) to obtain TEM00.
The alignment of the HWPs and the cube PBS didn't change the mode. Thermal lense of the cube?
- I could not lock the MC with the incident power below 100mW. So the BS in the MC REFL path was replaced by a total reflector (Y1-45S).
- This increased the power on the MC REFL PD x10 of the previous. NOW WE ARE CONSTRAINED BETWEEN 81deg~90deg. DON'T ROTATE FURTHER!
- The original BS was stored on the AP table as shown in the picture.
- This total reflector disabled the MC WFS QPDs. We can't use them.
[Lock of the MC with 20mW incident]
- Disable the MC autolocker.
- Disable the MC WFS.
- Run
/cvs/cds/caltech/scripts/MC/mcloopson
- Turn on the MCL servo.
- Set the MCL gain to 1.5 (it was nominally 0.3 for the high power)
- Just wait until lock.
[Gain boost after the lock] ...If you like to have more gain
- There was almost no room to increase the MCL gain.
- MC_REFL_GAIN can be increased from +6dB to +20dB
ezcawrite "C1:IOO-MC_REFL_GAIN" 20
- MC_VCO_GAIN can be increased from -3dB to +2dB
ezcawrite "C1:IOO-MC_VCO_GAIN" 2
- Crank the FSS gains
ezcawrite C1:PSL-FSS_MGAIN `ezcaread -n C1:PSL-STAT_FSS_NOM_C_GAIN`
ezcawrite C1:PSL-FSS_FASTGAIN `ezcaread -n C1:PSL-STAT_FSS_NOM_F_GAIN`
[If lock is lost]
- Run
/cvs/cds/caltech/scripts/MC/mcdown |
Attachment 1: MC_incident.png
|
|
Attachment 2: MC_REFL.png
|
|
2629
|
Mon Feb 22 21:07:26 2010 |
Jenne | Update | COC | Magnets glued to ITMX |
[Kiwamu, Jenne]
The magnets + dumbbell standoffs have been glued to ITMX. We're waiting overnight for them to dry.
Since I broke one of the magnet + dumbbells on the ITMY set, we've glued another dumbbell to the 6th magnet, and it should be ready for us to glue to ITMY tomorrow, once ITMX is dry and out of the fixture. This doesn't put us behind schedule at all, so that's good.
We had been concerned that there might be a problem with the arm of the guiderod fixture being glued to ITMY, but it was fine after all. Everything is going smoothly so far.
[Zach, Mott]
Zach and Mott are almost prepared to start cutting the viton for the earthquake stops. We need 2 full sets by Wednesday morning, when we expect to begin hanging the ITMs. |
2628
|
Mon Feb 22 13:08:27 2010 |
josephb | Update | Computers | Minor tweaks to c1omc |
While working on c1omc, I created a .cshrc file in the controls home directory, and had it source the cshrc.40m file so that useful shortcuts like "target" and "c" work, among other things. I also fixed the resolv.conf file so that it correctly uses linux1 as its name server (speeding up ssh login times). |
2627
|
Mon Feb 22 12:48:31 2010 |
josephb, alex, koji | Update | Computers | FE machines now coming up |
Even after bringing up the fb40m, I was unable to get the front ends to come up, as they would error out with an RFM problem.
We proceeded to reboot everything I could get my hands on, although its likely it was daqawg and daqctrl which were the issue, as on the C0DAQ_DETAIL screen their status had been showing as 0xbad, but after the reboot showed up as 0x0. They had originally come up before the frame builder had been fixed, so this might have been the culprit. In the course of rebooting, I also found c1omc and c1lsc had been turned off as well, and turned them on.
After this set of reboots, we're now able to bring the front ends up one by one. |
2626
|
Mon Feb 22 11:46:55 2010 |
josephb | Update | Computers | fb40m |
I fixed the JetStor 416S raid array IP address by plugging in my laptop to its ethernet port, setting my IP to be on the same subnet, and using the web interface. (After finally tracking down the password, it has been placed in the usual place).
After this change, I powered up the fb40m2 machine and reboot the fb40m machine. This seems to have made all the associated lights green.
Data viewer is working such that is recording from the point I fixed the JetStor raid array and did the fb40m reboot. It also can go back in time before the IP switch over. |
2625
|
Mon Feb 22 11:42:48 2010 |
Koji | Update | General | Prep for Power Supply Stop |
Turned on the power supply for the oplev lasers.
Turned on the power of the aux NPRO.
Turned on some of the Sorensen at 1X1.
Fixed the thermal output to round -4.0.
Locked PMC / MZ.
Waiting for the computers recovering. |
2624
|
Mon Feb 22 11:38:05 2010 |
joe, jenne and steve | Configuration | VAC | vacuum is back to normal |
Morning condition: vacuum rack power is still off, no MEDM screen reading.....meaning unknown vacuum pressure.We closed PSL shutter immediately.
Joe restored c1iscepis and Jenne powered up the vac-rack UPS. Now the rest of the vac-rack power were restored from starting at the top to bottom.
P1 was reading 15 mTorr. We restarted pumps and set vacuum valve positions. V1 opening required Rob's recipe of elog # 1863 to defeat interlock that
has a non communicating gauge: PTP1
CC1 pressure just reached 1e-6 Torr at VAC NORMAL configuration. |
2623
|
Mon Feb 22 10:25:37 2010 |
Jenne | Update | COC | ITMY standoff and guiderod epoxied |
This work happened on Friday, after Nodus and the elog went down....
[Jenne, Kiwamu]
The guiderod and standoff for ITMY were epoxied, and left drying over the weekend on the flow bench under a foil tent. The flow bench was off for the weekend, so we made tents which hopefully didn't have any place for dust to get in and settle on the mirrors.
There is a small chance that there will be a problem with glue on the arm of the fixture holding the guiderod to the optic. Kiwamu and I examined it, and hopefully it won't stick. We'll check it out this afternoon when we start getting ready for gluing magnets onto optics this afternoon. |
2622
|
Mon Feb 22 09:45:34 2010 |
josephb | Update | General | Prep for Power Supply Stop |
Quote: |
Autoburts have not been working since the network changeover last Thursday.
Last snapshot was around noon on Feb 11... 
It turns out this happened when the IP address got switched from 131.... to 192.... Here's the horrible little piece of perl code which was failing:
$command = "/usr/sbin/ifconfig -a > $temp";
system($command);
open(TEMP,$temp) || die "Cannot open file $temp\n";
$site = "undefined";
#
# this is a horrible way to determine site location
while ($line = <TEMP>) {
if ($line =~ /10\.1\./) {
$site = "lho";
} elsif ($line =~ /10\.100\./) {
$site = "llo";
} elsif ($line =~ /192\.168\./) {
$site = "40m";
}
}
if ($site eq "undefined") {
die "Cannot Determine Which LIGO Observatory this is\n";
I've now put in the correct numbers for the 40m...and its now working as before. I also re-remembered how the autoburt works:
1) op340m has a line in its crontab to run /cvs/cds/caltech/burt/autoburt/burt.cron (I've changed this to now run at 7 minutes after the hour instead of at the start of the hour).
2) burt.cron runs /cvs/cds/scripts/autoburt.pl (it was using a perl from 1999 to run this - I've now changed it to use the perl 5.8 from 2002 which was already in the path).
3) autoburt.pl looks through every directory in 'target' and tries to do a burt of its .req file.
Oh, and it looks like Joe has fixed the bug where only op440m could ssh into op340m by editing the host.allow or host.deny file (+1 point for Joe).
But he forgot to elog it (-1 point for Joe).®
|
I knew there was going to be a script somewhere with a hard coded IP address. My fault for missing it. However, in regards to the removal of op340m's host.deny file, I did elog it here. Item number 5. |
2621
|
Mon Feb 22 07:25:58 2010 |
rana | Update | General | Prep for Power Supply Stop |
Autoburts have not been working since the network changeover last Thursday.
Last snapshot was around noon on Feb 11... 
It turns out this happened when the IP address got switched from 131.... to 192.... Here's the horrible little piece of perl code which was failing:
$command = "/usr/sbin/ifconfig -a > $temp";
system($command);
open(TEMP,$temp) || die "Cannot open file $temp\n";
$site = "undefined";
#
# this is a horrible way to determine site location
while ($line = <TEMP>) {
if ($line =~ /10\.1\./) {
$site = "lho";
} elsif ($line =~ /10\.100\./) {
$site = "llo";
} elsif ($line =~ /192\.168\./) {
$site = "40m";
}
}
if ($site eq "undefined") {
die "Cannot Determine Which LIGO Observatory this is\n";
I've now put in the correct numbers for the 40m...and its now working as before. I also re-remembered how the autoburt works:
1) op340m has a line in its crontab to run /cvs/cds/caltech/burt/autoburt/burt.cron (I've changed this to now run at 7 minutes after the hour instead of at the start of the hour).
2) burt.cron runs /cvs/cds/scripts/autoburt.pl (it was using a perl from 1999 to run this - I've now changed it to use the perl 5.8 from 2002 which was already in the path).
3) autoburt.pl looks through every directory in 'target' and tries to do a burt of its .req file.
Oh, and it looks like Joe has fixed the bug where only op440m could ssh into op340m by editing the host.allow or host.deny file (+1 point for Joe).
But he forgot to elog it (-1 point for Joe).®
|
2620
|
Sun Feb 21 17:44:35 2010 |
rana | Update | General | Prep for Power Supply Stop |
- Turned on the RAID attached to linux1 (its our /cvs/cds disk)
- Turned on linux1 (it needed a keyboard and monitor in order to be happy - no fsck required)
- Turned on nodus (and started ELOG) + all the control room machines
- Turned on B/W monitors
- Untaped fridge
- Found several things OFF which were not listed in the Wiki...
- Turned ON the 2 big isolation transformers (next to Steve's desk and under the printer). These supply all of the CDS racks inside.
- ~75% of the power strips were OFF in the CDS racks ?? I turned on as many as I could find (except the OMC).
- Switched on and keyed on all of the FE and SLOW crates in no particular order. Some of the fans sound bad, but otherwise OK.
- Turned on all of the Sorensens that are labeled.
- Turned ON the linear supplies close to the LSC rack.
- ON the Marconis - set them according to the labels on them (probably out-dated).
- After restoring power to the PSL enclosure (via the Isolation Transformer under the printer) turned the Variac ON and HEPA on full speed.
- Plugged in the PSs for the video quads. Restored the Video MUX settings - looks like we forgot to save the correct settings for this guy...
PSL
1) Turned on the chiller, then the MOPA, then the RC's Heater power supply.
2) Shutter is open, laser is lasing, PMC is locked.
3) RC temperature is slowly rising. Will probably be thermalized by tomorrow.
Sun Feb 21 20:04:17 2010
Framebuilder is not mounting its RAID frames - in fact, it doesn't mount anything because the mountall command is failing on the RAID with the frames. The Jetstor RAID is also not responding to ping. Looks like the JetStor RAID which has all of our frames is still on the old 131 network, Joe. |
2619
|
Fri Feb 19 16:40:43 2010 |
kiwamu | Update | Green Locking | rearrange the optics on the end table |
Koji and kiwamu
The existing optics on the ETMX/ETMY end table were rearranged in this morning.
The main things we have done are -
1. relocation of the optical levers for ETMs ( as mentioned in koji's entry )
This relocation can make a space so that we can setup the green locking stuffs.
The optical path of the green locking is planed to start from the right top corner on the table, therefore we had to relocate the oplevs toward the center of the table.
2. relocation of the lens just before the tube
Because we are going to shoot the green beam into the arm cavity, we don't want to have any undesired lenses before the cavity.
For this reason we changed the position of the lens, it was standing just in front of the tube, now it's standing on the left side of the big mirror standing center top.
Since we did not find a significant change in its the spot size of the transmitted light, we did not change the position of all the TRANS_MON_PDs and its mirrors. And they are now well aligned.
Attachment1: ETMX end table
Attachment2: ETMY end table |
Attachment 1: DSC_1202.JPG
|
|
Attachment 2: DSC_1207.JPG
|
|